When 22.00 mL of 0.5000 M H_2SO_4 is added to 22.0
Answer #1

Enthalpy dH = q = mCpdT

with,

m = 44 g

Cp = 4.184 J/g.oC

dT = 30.17 - 23.50 = 6.67 oC

we get,

dH = 44 x 4.184 x 6.67 = 1.228 kJ

dH per mole of H2SO4,

dH = 1.228/0.5 M x 0.022 ml = 111.636 kJ/mol